# Periodicity of traces of Hecke operators modulo prime powers

### v1

@@ 3.1 The case of $\ell\geq 3$ and $\ell\nmid q$
Let $K$ be a field. Suppose that $\zeta\in K^{\times}$ is an element of order $m<\infty$. Then for all $A,B\in\overline{K}$, we have
@@
Let $\ell\geq 3$ be prime and assume $\ell\nmid q$. Let $s\geq 1$ and $1\leq t\leq s$. Put $m=\ell^{t-1}(\ell-1)/2$. Then the polynomial $d_{q,t}:=(1+qx^{2})^{2m}-x^{2m}$ divides $x^{n_{U}}-1$ modulo $\ell^{s+1-t}$.
@@
Finally, since $q$ is square, two of the $f_{i}$ are squares of linear factors. Therefore $D_{q}$ divides $(x^{\ell^{2}-1}-1)^{\ell}=x^{\ell(\ell^{2}-1)/2}-1$ mod $\ell$, which is what we wanted to show. ∎
@@
Recall that $\mathrm{eis}_{H,k}(q)=\allowbreak\mathrm{eis}_{H,k+2i}(q)$ for any $i\geq 0$, by Theorem 2.1. The result then follows from Lemmas 3.3, 3.6 and 3.17 after noting that $n_{N}\mid n_{U}$ and that $2\mid n_{U}$. ∎
@@ 3.2 The case of $\ell=2$ and $\ell\nmid q$
We have that $d_{q,1}=1+qx^{2}-x$ divides $x^{\ell^{2}-1}-1$ mod $\ell$ and hence that $d_{q,t}\equiv d_{q,1}^{\ell^{t-1}}$ divides $(x^{\ell^{2}-1}-1)^{\ell^{t-1}}=x^{\ell^{t-1}(\ell^{2}-1)}-1$. We conclude by applying Lemma 3.13 with r=str=s-t. ∎
@@
If $t\geq 2$, this follows in the same way as Lemma 3.16. For $t=1$ we have by definition of $f_{r,m,k}$ that
@@
Let $\ell=2$ be prime, $s\geq 1$ and suppose $\ell\nmid q$. For any $k\geq 0$, we have
@@
$\equiv_{\ell^{s}}f_{0,1,k}\,b_{2r+\delta_{k},1}+\sum_{t=2}^{s}\ell^{t-1}\sum_{r=0}^{m_{\ell,t}-1}f_{r,m_{\ell,t},k}\,b_{2r+\delta_{k},t},$
@@
Recall that $\mathrm{eis}_{H,k}(q)=\mathrm{eis}_{H,k+2i}(q)$ for any $i\geq 0$, by Theorem 2.1. The result then follows from Lemmas 3.3, 3.6 and 3.22. ∎
@@ 3.4 The case when $H$ is not representable and $\ell=2$ or 33
for all $0\leq j\leq\lfloor(\tilde{s}-1-\delta_{k})/2\rfloor$ and $(E,\varphi)]\in\mathcal{N}(\mathbb{F}_{q})$. After dividing this with $\operatorname{Aut}(E,\varphi)$ and summing over all $(E,\varphi)]\in\mathcal{N}(\mathbb{F}_{q})$ it follows that
@@
After dividing this with $\operatorname{Aut}(E,\varphi)$ and summing over all $(E,\varphi)]\in\mathcal{N}(\mathbb{F}_{q})$ it follows that
$U(k+2)-U(k+2+n)\equiv_{\ell^{s}}0$
@@
Finally, if $\ell=2$ and $\mathrm{Id}\in H$, then $1\in\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)$ for any $(E,\varphi)\in\mathcal{Y}(\mathbb{F}_{q})$. Twisting $(E,\varphi)\in\mathcal{Y}(\mathbb{F}_{q})$ by the automorphism 1-1 gives an element $(E^{(-1)},\varphi)\in\mathcal{Y}(\mathbb{F}_{q})$ with $a_{1}(E^{(-1)})=-a_{1}(E)$. If we fix $t\neq 0$ then sending $E$ to $E^{(-1)}$ induces a bijection between the elements $(E,\varphi)]\in[\mathcal{Y}(\mathbb{F}_{q})]$ with $a_{1}(E)=t$ and the elements $(E^{\prime},\varphi^{\prime})]\in[\mathcal{Y}(\mathbb{F}_{q})]$ such that $a_{1}(E^{\prime})=-t$. This gives an extra divisibility by $2$ when going through the arguments above. ∎
@@ 4 Proof of the main theorem for Drinfeld modular forms
  • Put $A=\mathcal{O}_{C}(C\setminus\{\infty\})$, for a smooth geometrically connected curve $C/\mathbb{F}_{q}$, where $\infty$ is a point of degree 1.
@@ 4.1 The case of $\mathfrak{l}\neq\mathfrak{p}$
Let $f\in A[x]$ and $t\in\mathbb{Z}_{\geq 0}$. Suppose that $f$ divides $x^{n}-1$ mod $\mathfrak{l}$. Then $f^{p^{t}}$ divides $x^{p^{t}n}-1$ mod $\mathfrak{l}^{p^{t}}$.
@@
$n_{U}=\begin{cases}\text{lcm}(2,p^{\tilde{s}}(|\mathfrak{l}|^{2}-1))&\text{if }p=2;\\ p^{\tilde{s}}(|\mathfrak{l}|^{2}-1)&\text{if }p>2,\ \deg(\mathfrak{l})\equiv_{2}1\text{ and }\left(\frac{\wp_{n}}{\mathfrak{l}}\right)=-1;\\ p^{1+\tilde{s}}(|\mathfrak{l}|^{2}-1)/2&\text{if }p>2,\ \deg(\mathfrak{l})\equiv_{2}1\text{ and }\left(\frac{\wp_{n}}{\mathfrak{l}}\right)=1;\\ p^{1+\tilde{s}}(|\mathfrak{l}|^{2}-1)&\text{otherwise.}\end{cases}$
@@
The period of the trace is $L:=\text{lcm}(n_{N},n_{U})$. If $p=2$, then $L=2p^{\tilde{s}}(|\mathfrak{l}|^{2}-1)=p^{1+\tilde{s}}(|\mathfrak{l}|^{2}-1)$. If $p\neq 2$, note that if $\wp_{n}$ is a square mod $\mathfrak{l}$ and $\deg(\mathfrak{l})$ is even, then $\lambda\wp_{n}$ is a square mod $\mathfrak{l}$ for all $\lambda\in\mathbb{F}_{q}^{\times}$. Hence $n_{N}\mid n_{U}$ in all cases, so $L=n_{U}$. ∎
@@ 4.2 The case of $\mathfrak{l}=\mathfrak{p}$
for all $k\geq 2\lfloor(s-1)/n\rfloor$, $m\geq 1$, $b\in\mathbb{F}_{q}^{\times}$ and $r$.
@@ 4.3 Traces modulo $\infty$
Then $\operatorname{Tr}_{\infty}(\mathbf{T}^{n}_{\mathfrak{p}}\,|\,\operatorname{S}_{k+2,l}({\mathcal{K}}))\in\mathcal{O}_{\infty}$; this follows from Theorem 2.8 and the Riemann hypothesis for Drinfeld modules over finite fields, which says that $2v_{\infty}(a_{\varphi})\leq-v_{\infty}(\wp_{n})$, see [gek_finite, Theorem 5.1]. If $A=\mathbb{F}_{q}[T]$, then $v_{\infty}=-\deg_{T}$ and the strong Ramanujan bound [devries_traces, Conjecture 4.3] is equivalent to the statement
$\operatorname{Tr}_{\infty}(\mathbf{T}^{n}_{\mathfrak{p}}\,|\,\operatorname{S}_{k+2,l}(\mathrm{GL}_{2}(A)))\equiv 0\pmod{\pi_{\infty}^{n\deg(\mathfrak{p})(q-1)/2}}\qquad\text{for all }k,l.$
@@
$\operatorname{Tr}_{\infty}(\mathbf{T}_{\mathfrak{p}}^{n}\,|\,\operatorname{S}_{k+2,l}({\mathcal{K}}))\equiv_{\pi_{\infty}^{s}}N_{\infty}(k+2,l)+U_{\infty}(k+2,l),$
@@ 5.1 Elliptic modular forms of level $1$ for small $\ell^{s}$
$\sum_{[(E,\varphi)]\in\mathcal{[}Y_{H}(\mathbb{F}_{q})]}\frac{h_{i}(a_{1}(E))}{\#\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)}\equiv_{\ell^{t_{i}}}[a_{1}^{k}]+\sum_{j=1}^{i}c_{i,j}\,[a_{1}^{k-j}]\equiv_{\ell^{t_{i}}}0.$ (8)
@@
Using the results of Section 5.1 together with [silverman, Theorem 10.1], we find that
@@
For results with $\ell\leq 7$, $s=1$ and $\ell\neq p$, see [CFW, Theorem 2a].
@@
where $H(\Delta)$ denotes the Kronecker class number of $\Delta$, see [Schoof, Theorem 4.6].
@@ 5.2 Congruence equations over function fields
In the setting of the previous example, the Drinfeld modular of lowest weight which does not possess an $A$-expansion is $)gh^{2}\in\operatorname{S}_{3q+1,2}(\mathrm{GL}_{2}(\hat{A}$)). One can use this to show that the elements $c_{k,l}]$ from Theorem 2.8 lie in $\mathbb{F}_{p}\subset\mathbb{F}_{q}[T]$ for k<3q1k<3q-1.
In particular, if $q=3$ and $\mathfrak{l}$ is a prime of degree 1, we have $a^{k+6}\equiv_{\mathfrak{l}^{2}}a^{k}$ for all $k\geq 2$ and all $a\in A$. It then follows from Theorem 2.8 that for all $k$ and $l$, $\operatorname{Tr}(\mathbf{T}_{\mathfrak{p}}^{n}\,|\,\operatorname{S}_{k+2,l}(\mathrm{GL}_{2}(\widehat{\mathbb{F}_{3}[T]}))\pmod{\mathfrak{l}^{2}}$ is a polynomial in $\wp^{n}$ over $\mathbb{F}_{3}$. In particular, this trace depends only on $\wp^{n}\pmod{\mathfrak{l}^{2}}$, analogous to Example 5.1.
@@
Using Theorem 1.4, this could be verified in a finite number of computations.
@@ 5.3 Hecke polynomials
If $H$ is not representable, it is in some cases still possible to deduce the conclusion of the proposition by applying Newton’s identities to the traces of $T(p)$ on the spaces $W_{k}(H)=\widetilde{S}_{k+p-1}(H)/E_{p-1}\widetilde{S}_{k}(H)$, where $\widetilde{S}_{k}(H)$ is the reduction modulo $p$ of $S_{k}(H)$ and $E_{p-1}$ denotes the Eisenstein series of weight p1p-1, which satisfies $E_{p-1}\equiv_{p}1$. Since Newton’s identities involve denominators, one needs to assume that $p>\dim_{\mathbb{F}_{p}}W_{k}(H)$; this holds for instance if $H$ corresponds to $\Gamma_{1}(N)$ with $N\leq 3$ (and $\Gamma_{1}(N)$ is representable for $N\geq 4$, see [KatzMazur, Cor. 2.7.3]).
The same proof works for Drinfeld modular forms, which gives the following result. Note that $\mathcal{K}$ is representable if it is admissible in the sense of [bockle].
@@
for all $k\geq 3$ and any $l\in\mathbb{Z}$.
Define a $p$-adic slope of $T(p)$ to be the normalised $p$-adic valuation of a root of $\mathfrak{f}_{k}(x)\in\mathbb{Z}_{p}[x]$, and denote by $d_{p}(k,\alpha)$ its multiplicity. Similarly, for Drinfeld modular forms, denote by $d_{\mathfrak{p}}(k,l,\alpha)$ the multiplicity of the $\mathfrak{p}$-adic slope $\alpha$ on $\operatorname{S}_{k,l}(\mathcal{K}))$.
@@
In particular, with the exception of $k=2$, we recover Hida’s result for elliptic modular forms that the Gouvêa–Mazur conjecture holds for slope 0.

### v2

@@ 3.1 The case of $\ell\geq 3$ and $\ell\nmid q$
Let $K$ be a field. Suppose that $\zeta\in K^{\times}$ is an element of order $m<\infty$. Then for all $A,B\in\bar{K}$, we have
@@
Let $\ell\geq 3$ be prime and assume $\ell\nmid q$. Let $s\geq 1$ and $1\leq t\leq s$. Put $m=m_{\ell,t}=\ell^{t-1}(\ell-1)/2$. Then the polynomial $d_{q,t}:=(1+qx^{2})^{2m}-x^{2m}$ divides $x^{n_{U}}-1$ modulo $\ell^{s+1-t}$.
@@
Finally, since $q$ is square, two of the $f_{i}$ are squares of linear factors. Therefore $D_{q}$ divides $(x^{(\ell^{2}-1)/2}-1)^{\ell}=x^{\ell(\ell^{2}-1)/2}-1$ mod $\ell$, which is what we wanted to show. ∎
@@
Recall that $\mathrm{eis}_{H,k}(q)=\mathrm{eis}_{H,k+2i}(q)$ for any $i\geq 0$, by Theorem 2.1. Together with Lemmas 3.3, 3.6 and 3.17, this gives the period $\text{lcm}(2,n_{N},n_{U})=n_{U}$, where $n_{N}\mid n_{U}$ because $\ell^{2}-1\equiv_{8}0$. ∎
@@ 3.2 The case of $\ell=2$ and $\ell\nmid q$
We have that $d_{q,1}=1+qx^{2}-x$ divides $x^{\ell^{2}-1}-1$ mod $\ell$ and hence that $d_{q,t}\equiv_{\ell}d_{q,1}^{\ell^{t-1}}$ divides $(x^{\ell^{2}-1}-1)^{\ell^{t-1}}=x^{\ell^{t-1}(\ell^{2}-1)}-1$. We conclude by applying Lemma 3.13 with r=str=s-t. ∎
@@
If $t\geq 2$, this follows in the same way as Lemma 3.16. For $t=1$, the definition of $f_{r,m,k}$ gives
@@
Let $\ell=2$, $s\geq 1$ and suppose $\ell\nmid q$. For any $k\geq 0$, we have
@@
$\equiv_{\ell^{s}}f_{0,1,k}\,b_{\delta_{k},1}+\sum_{t=2}^{s}\ell^{t-1}\sum_{r=0}^{m_{\ell,t}-1}f_{r,m_{\ell,t},k}\,b_{2r+\delta_{k},t},$
@@
Recall that $\mathrm{eis}_{H,k}(q)=\mathrm{eis}_{H,k+2i}(q)$ for any $i\geq 0$, by Theorem 2.1. By Lemmas 3.3, 3.6 and 3.22, the period is $\text{lcm}(2,n_{N},n^{\prime}_{U})$, which gives the result. ∎
@@ 3.4 The case when $H$ is not representable and $\ell=2$ or 33
for all $0\leq j\leq\lfloor(\tilde{s}-1-\delta_{k})/2\rfloor$ and $(E,\varphi)]\in\mathcal{N}(\mathbb{F}_{q})$. After dividing this by $\operatorname{Aut}(E,\varphi)$ and summing over all $(E,\varphi)]\in\mathcal{N}(\mathbb{F}_{q})$ it follows that
@@
After dividing this by $\operatorname{Aut}(E,\varphi)$ and summing over all $(E,\varphi)]\in\mathcal{N}(\mathbb{F}_{q})$ it follows that
$U(k+2)-U(k+2+n)\equiv_{\ell^{s}}0.$
@@
If $\ell=2$ and $\mathrm{Id}\in H$, then $1\in\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)$ for any $(E,\varphi)\in\mathcal{Y}_{H}(\mathbb{F}_{q})$. Twisting $(E,\varphi)\in\mathcal{Y}_{H}(\mathbb{F}_{q})$ by the automorphism 1-1 gives an element $(E^{(-1)},\varphi)\in\mathcal{Y}_{H}(\mathbb{F}_{q})$ with $a_{1}(E^{(-1)})=-a_{1}(E)$. If we fix $t\neq 0$ then sending $E$ to $E^{(-1)}$ induces a bijection between the elements $(E,\varphi)]\in[\mathcal{Y}_{H}(\mathbb{F}_{q})]$ with $a_{1}(E)=t$ and the elements $(E^{\prime},\varphi^{\prime})]\in[\mathcal{Y}_{H}(\mathbb{F}_{q})]$ such that $a_{1}(E^{\prime})=-t$. This gives an extra divisibility by $2$ when going through the arguments above.
Finally, in the case $\ell=p$, the arguments from Section 3.3 work also in characteristic $2$ (respectively $3$), with $s$ replaced by $s+\nu_{2}(H,q)$ or $s+\nu_{2}(H,q)-1$ if $\mathrm{Id}\in H$ (respectively $s+\nu_{3}(H,q)$). ∎
@@ 4 Proof of the main theorem for Drinfeld modular forms
  • Put $A=\mathcal{O}_{C}(C\setminus\{\infty\})$ for a smooth geometrically connected curve $C/\mathbb{F}_{q}$, where $\infty$ is a point of degree 1.
@@ 4.1 The case of $\mathfrak{l}\neq\mathfrak{p}$
Let $f\in A[x]$ and $t\geq 0$. Suppose that $f$ divides $x^{n}-1$ mod $\mathfrak{l}$. Then $f^{p^{t}}$ divides $x^{p^{t}n}-1$ mod $\mathfrak{l}^{p^{t}}$.
@@
$n_{U}=\begin{cases}\text{lcm}(2,p^{\tilde{s}}(|\mathfrak{l}|^{2}-1))&\text{if }p=2;\\ p^{\tilde{s}}(|\mathfrak{l}|^{2}-1)&\text{if }p>2,\ \deg(\mathfrak{l})\equiv_{2}0\text{ and }\left(\frac{\wp_{n}}{\mathfrak{l}}\right)=-1;\\ p^{\tilde{s}+1}(|\mathfrak{l}|^{2}-1)/2&\text{if }p>2,\ \deg(\mathfrak{l})\equiv_{2}0\text{ and }\left(\frac{\wp_{n}}{\mathfrak{l}}\right)=1;\\ p^{\tilde{s}+1}(|\mathfrak{l}|^{2}-1)&\text{otherwise.}\end{cases}$
@@
The period of the trace is $L:=\text{lcm}(n_{N},n_{U})$. If $p=2$, then $L=2p^{\tilde{s}}(|\mathfrak{l}|^{2}-1)=p^{\tilde{s}+1}(|\mathfrak{l}|^{2}-1)$. If $p\neq 2$, note that if $\wp_{n}$ is a square mod $\mathfrak{l}$ and $\deg(\mathfrak{l})$ is even, then $\lambda\wp_{n}$ is a square mod $\mathfrak{l}$ for all $\lambda\in\mathbb{F}_{q}^{\times}$. Hence $n_{N}\mid n_{U}$ in all cases, so $L=n_{U}$. ∎
@@ 4.2 The case of $\mathfrak{l}=\mathfrak{p}$
for all $k\geq 2\lfloor(s-1)/n\rfloor$, $m\geq 1$, $b\in\mathbb{F}_{q}^{\times}$ and $r\in\mathbb{Z}$.
@@ 4.3 Traces modulo $\infty$
Then $\operatorname{Tr}_{\infty}(\mathbf{T}^{n}_{\mathfrak{p}}\,|\,\operatorname{S}_{k+2,l}({\mathcal{K}}))\in\mathcal{O}_{\infty}$; this follows from Theorem 2.8 and the Riemann hypothesis for Drinfeld modules over finite fields, which says that $2v_{\infty}(a_{\varphi})\leq-v_{\infty}(\wp_{n})$, see [8, Theorem 5.1]. If $A=\mathbb{F}_{q}[T]$, then $v_{\infty}=-\deg_{T}$ and the strong Ramanujan bound [5, Conjecture 4.3] is equivalent to the statement
$\operatorname{Tr}_{\infty}(\mathbf{T}^{n}_{\mathfrak{p}}\,|\,\operatorname{S}_{k+2,l}(\mathrm{GL}_{2}(A)))\equiv 0\pmod{\pi_{\infty}^{n\deg(\mathfrak{p})(q-1)/2}}\qquad\text{for all }k,l\in\mathbb{Z}.$
@@
$\operatorname{Tr}_{\infty}(\mathbf{T}_{\mathfrak{p}}^{n}\,|\,\operatorname{S}_{k+2,l}({\mathcal{K}}))\equiv_{\pi_{\infty}^{s}}-N_{\infty}(k+2,l)-U_{\infty}(k+2,l),$
@@ 5.1 Elliptic modular forms of level $1$ for small $\ell^{s}$
$\sum_{[(E,\varphi)]\in[\mathcal{Y}_{H}(\mathbb{F}_{q})]}\frac{h_{i}(a_{1}(E))}{\#\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)}\equiv_{\ell^{t_{i}}}[a_{1}^{k}]+\sum_{j=1}^{i}c_{i,j}\,[a_{1}^{k-j}]\equiv_{\ell^{t_{i}}}0.$ (8)
@@
Using the results of Section 5.1 together with [19, Theorem 10.1], we find that
@@
For results with $\ell\leq 7$, $s=1$ and $\ell\neq p$, see [3, Theorem 2a].
@@
where $H(\Delta)$ denotes the Kronecker class number of $\Delta$, see [17, Theorem 4.6].
@@ 5.2 Congruence equations over function fields
In the setting of the previous example, the Drinfeld cusp form of lowest weight which does not possess an $A$-expansion is $)gh^{2}\in\operatorname{S}_{3q+1,2}(\mathrm{GL}_{2}(\hat{A}$)). One can use this to show that the elements $c_{k,l}]$ from Theorem 2.8 lie in $\mathbb{F}_{p}\subset\mathbb{F}_{q}[T]$ for k<3q1k<3q-1.
In particular, if $q=3$ and $\mathfrak{l}$ is a prime of degree 1, we have $a^{k+6}\equiv_{\mathfrak{l}^{2}}a^{k}$ for all $k\geq 2$ and all $a\in A$. It then follows from Theorem 2.8 that for all $k$ and $l$, $\operatorname{Tr}(\mathbf{T}_{\mathfrak{p}}^{n}\,|\,\operatorname{S}_{k+2,l}(\mathrm{GL}_{2}(\mathbb{F}_{3}[T])))\pmod{\mathfrak{l}^{2}}$ is a polynomial in $\wp^{n}$ over $\mathbb{F}_{3}$. In particular, this trace depends only on $\wp^{n}\pmod{\mathfrak{l}^{2}}$, analogously to Example 5.1.
@@
Using Theorem 1.4, this could be verified with a finite number of computations.
@@ 5.3 Hecke polynomials
If $H$ is not representable, it is in some cases still possible to deduce the conclusion of the proposition by applying Newton’s identities to the traces of $T(p)$ on the spaces $W_{k}(H)=\widetilde{S}_{k+p-1}(H)/E_{p-1}\widetilde{S}_{k}(H)$, where $\widetilde{S}_{k}(H)$ is the reduction modulo $p$ of $S_{k}(H)$ and $E_{p-1}$ denotes the Eisenstein series of weight p1p-1, which satisfies $E_{p-1}\equiv_{p}1$. Since Newton’s identities involve denominators, one needs to assume that $p>\dim_{\mathbb{F}_{p}}W_{k}(H)$; this holds for instance if $H$ corresponds to $\Gamma_{1}(N)$ with $N\leq 3$ (and $\Gamma_{1}(N)$ is representable for $N\geq 4$, see [13, Cor. 2.7.3]).
The same proof works for Drinfeld modular forms, which gives the following result. Note that $\mathcal{K}$ is representable if it is admissible in the sense of [2].
@@
for any $k\geq 3$ and $l\in\mathbb{Z}$.
Define a $p$-adic slope of $T(p)$ to be the normalized $p$-adic valuation of a root of $\mathfrak{f}_{k}(x)\in\mathbb{Z}_{p}[x]$, and denote by $d_{p}(k,\alpha)$ its multiplicity. Similarly, for Drinfeld modular forms, denote by $d_{\mathfrak{p}}(k,l,\alpha)$ the multiplicity of the $\mathfrak{p}$-adic slope $\alpha$ on $\operatorname{S}_{k,l}(\mathcal{K})$. Retaining the representability assumption on the level, we deduce a dimension periodicity for slope 0.
@@
In particular, with the exception of $k=2$, this recovers Hida’s result for elliptic modular forms that the Gouvêa–Mazur conjecture holds for slope 0.